f four marbles are to be picked at random, in how many ways could we pick


One marble for each color.


All the green marbles


3 of the green marbles

1
Expert's answer
2021-12-01T18:17:39-0500

In a box, there are 3 red marbles,4 green marbles, 2 blue marbles and 5 yellow marbles


number of ways we could pick one marble for each color:

"n=3\cdot4 \cdot2\cdot5=120"


number of ways such that all the green marbles:

"n=C^4_4=1"


number of ways such that 3 are the green marbles:

"n=3+2+5=10" (number of not green marbles)